Brett's father is four times as old as Brett is now. In four years, Brett's age will beone-third of his father's age. How old are they now?
Answer by scott8148(6628)

f = 4b

f + 4 = 3(b + 4)

substituting ___ 4b + 4 = 3b + 12

b = 8

f = 32